OpenStudy (anonymous):

In a certain county, the number of charter schools is 10 less than twice the number of alternative schools. We know that there are 46 charter schools in the county. How many alternative schools are in the county?

OpenStudy (anonymous):

\[\let \ the \ number \ of \ alternative \ schools \ be \ x\\problem\ says, \ number \ of \ charter \ schools \ =2x-10\\again,\ given\ number\ of\ charter\ schools\ = 46\\ \implies \ 2x-10=46\\or\ 2x=46+10\\ or\ 2x= 56\\or\ x=\frac{ 56 }{ 2}\\ \therefore \ x= 28\\hence \ number \ of\ alternative\ schools\ are\ 28. Ans\]